OpenStudy (dubsubnation):

Solid sodium oxide reacts with excess water to produce 0.650 moles sodium hydroxide dissolved in water, as shown in the balanced chemical equation below. What mass (in grams) of sodium oxide was consumed? Na2O(s) + H2O(l) --> 2NaOH(aq) My answer is 25.4g, but that is wrong. The correct answer is 20.2 g. Can someone write out how to get that?

OpenStudy (lovemusic56):

\[.650 NaOH \times \frac{ 1 mol Na _{2}O }{ 2 mol NaOH} \times \frac{ 61.98 grams Na _{2}O }{ 1 mol Na _{2}O }\]